2. Solve the following multi-stage DDM

A stock is expected to pay 0.50, 0.55 and 0.60 euros per share in the next 3 years.
Then, a growth phase of 5 years at 5% is expected.
Finally, the future growth of the company is expected to be 2% per year.
This stock has a beta 1.1, the risk-free rate is 1%, the expected return of the market is 6%.
Calculate the fair value of this stock today.
Answer #1

required return=1%+1.1*(6%-1%)=6.5%
Fair value of the stock today=0.50/1.065+0.55/1.065^2+0.60/1.065^3+0.60/1.065^3*(1.05/1.065)+0.60/1.065^3*(1.05/1.065)^2+0.60/1.065^3*(1.05/1.065)^3+0.60/1.065^3*(1.05/1.065)^4+0.60/1.065^3*(1.05/1.065)^5+0.60/1.065^3*(1.05/1.065)^5*1.02/(6.5%-2%)=14.31956436